A right circular cone is inscribed in a cube having a diagonal which
measures 20√3 cm.
Altitude of the cone = 20 cm
Volume of the cone = 2094.4 cubic centimeter
Surface Area of the cone = 942.5 cm squared
1. What is the volume of the space between the cone and the cube?
1
Expert's answer
2020-11-25T15:36:21-0500

Firstly, we need to clarify the side of cube by it`s diagonal. if we consider a as a side of the cube,

then a2+2*(a)2=d2 this came from Pythagoras`s theorem.

In this case, a=20 cm.

Let`s continue the space between cube and cone:

Vspace=Vcube-Vcone,

Vcone is already given and we find Vcube =a3 through this form.

Let`s put above :

Vspace = a3-2094.4cm3=203-2094.4=5905.6 cm3.

Answer:5905.6 cm3.